OpenStudy (anonymous):

expand: log[10]:√4:(y^3)/(20x^5)

OpenStudy (kira_yamato):

????

OpenStudy (anonymous):

log[10]√4:(y^3)/(20x^5) =log[10]√4+log(10)(y^3)-log(10)(20x^5) =(1/2)log[10] 4 + 3log(10) y - 5log(10) 20